OpenStudy (anonymous):

find the derivative of the following functions y=ln (e^x)/(1+e^x)

OpenStudy (raden):

simplify the numerator, first. ln(e^x) = x ln(e) = x * 1 = x so, y = x/(1+e^x) then use the quotient rule : if y = u/v then y' = (vu' - uv')/v^2 try it ....

OpenStudy (anonymous):

According to what RadEn said , I think this is the answer \[\frac{ 1+e^x -xe^x }{ (1+e^x)^2}\]